Find the Least Common Multiple of the numbers or monomials.
23.8,18
Any help with this equation would be greatly appreciated:)

Answer by stanbon(75887)

Find the Least Common Multiple of the numbers or monomials.
23. 8,18
8 = 2^3
18 = 2*3^2
--------------
LCM contains each factor to its highest power
--------------
LCM = 2^3*3^2 = 8*9 = 72
==============
Cheers,
Stan H.